Owl shift 11/26/99

Sasha Glazov (glazov@hep.uchicago.edu)
Fri, 26 Nov 1999 07:20:50 -0600

Shift Summary for owl shift, 26 Nov 99.

On Shift: MT, SG

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 7.0
Hours of E799 data (special runs not included): 6h30min
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: debuging of hcc problems
kfrends
drt_fresh_start

========================================================================
Runs ended this shift:
Run 15057 good data. HCC problems for spills 215-266. Run ended due
to alarm on "L3 accept fraction"

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:
Realtively smooth 799 datataking. HCC problems for ~50 spills

> ========================================================================
Description of Shift:



run spill time comment
------------------------------------------------------------------------
15057 0:00 Start of the shift
215 3:30 Alarm on L2/L1 fraction. After some time found
that the reason is hot HCC bit #701.
3:55 Paged Jason
4:25 Jason is here. First he tries to swap cables ->
255-261 HCC cable swap
266 4:40 After Jason plugs the cable back in the hot
bit seems to disapear. L2/L1 is back to normal.

4:40 Scaler HA2/SEm exeeds upper limit. Checked online
plots for HA2, all looks fine --> the bad scaler
problem ?

6:00 Run wants to be stoped: "Please check L3 accept
fraction. Do drt_fresh_start

15058 6:15 Start new run...

6:50 Tapes are full ... We the meter was only 73% full

========================================================================
Follow up Notes on Previous Shift

Done with drt_fresh_start
========================================================================
Follow up notes to next shift: